OpenStudy (anonymous):

On the Planet X the height of a rock thrown up is h(t) = 32t - 4t^2 meters in seconds. a. What is the equation for the velocity and acceleration of the rock? b. What is the initial velocity? c. How long will it take the rock to reach its maximum height? d. What is the max. height? e. What is the acceleration at t-=10 seconds?

jimthompson5910 (jim_thompson5910):

I'm assuming this if for a calculus class?

jimthompson5910 (jim_thompson5910):

If so, then a) Position: h(t) = 32t - 4t^2 Velocity: v(t) = h'(t) = 32 - 8t Acceleration: a(t) = v'(t) = -8

jimthompson5910 (jim_thompson5910):

b) Initial velocity is when t = 0, so v(t) = 32 - 8t v(0) = 32 - 8(0) v(0) = ???

OpenStudy (anonymous):

Yes CALC

OpenStudy (anonymous):

v= 32

jimthompson5910 (jim_thompson5910):

c) Set the velocity function equal to zero and solve for t v(t) = 32 - 8t 0 = 32 - 8t 32 - 8t = 0 t = ??

jimthompson5910 (jim_thompson5910):

good, so the initial velocity is 32 m/s

OpenStudy (anonymous):

Is initial velocity 0?

jimthompson5910 (jim_thompson5910):

no, 32 m/s

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

How long...?

jimthompson5910 (jim_thompson5910):

d. What is the max. height? Take the answer you get from part c) and plug it into h(t) = 32t - 4t^2

jimthompson5910 (jim_thompson5910):

when 32 - 8t = 0, t = ???

jimthompson5910 (jim_thompson5910):

e. What is the acceleration at t = 10 seconds the acceleration function is a(t) = -8 and this is found in part a this is a constant function, so it will NEVER change the acceleration therefore is ALWAYS -8 m/s^2

jimthompson5910 (jim_thompson5910):

so on this unknown planet, objects coming closer to it will speed up by 8 m/s every second they get closer to the planet

jimthompson5910 (jim_thompson5910):

b. What is the initial velocity? 32 m/s (shown above)

OpenStudy (anonymous):

Awesome! Thanks a ton!

OpenStudy (anonymous):

How do I give you points or whatever??

jimthompson5910 (jim_thompson5910):

you just did, thx